OpenStudy (anonymous):

How to solve this integral:

OpenStudy (anonymous):

\[\int\limits_{}^{}1/(1-x ^{2})\]

OpenStudy (anonymous):

what does that long thing before the first one mean?

OpenStudy (anonymous):

hint:x=sinu....dx=cosu.du

OpenStudy (anonymous):

ln(x+1/1-x^2)

OpenStudy (anonymous):

i still don't get it

OpenStudy (anonymous):

SKforlife: It's the symbol for integral equations :D

OpenStudy (anonymous):

integ_(cosu.du/1-sin^2(u))=integ_cosudu/cos^2(u)=integ_du/cosu=integ_secu.du integ_secu.du=ln(secu+tanu) draw a triangle..your angle is u..sinu=x then write each other.

OpenStudy (lalaly):

i\[=\int\limits{\frac{1}{(1-x)(1+x)}dx}\]break intp partial fractions\[\frac{1}{(1-x)(1+x)}=\frac{A}{1-x}+\frac{B}{1+x}\]

OpenStudy (nikvist):

\[\frac{1}{1-x^2}=\frac{1}{2}\left(\frac{1}{1-x}+\frac{1}{1+x}\right)\]

OpenStudy (lalaly):

yeah what nikvist did xD

OpenStudy (lalaly):

Did u understnd?

OpenStudy (anonymous):

yes, thanks

OpenStudy (lalaly):

:)